OpenStudy (anonymous):

In the book chapter6.2, it says when A and B both diagonalizable, AB=BA is true if and only if they have the same eigenvectors. It's easy to obtain that AB=BA,but how to prove vise versa??

OpenStudy (anonymous):

The statement does not depend of basis choice, so let's consider the basis, consisting from eigenvectors of A. Matrix presentation of A is diagonal in that basis. AB=BA implies that matrix presentation of B is a diagonal matrix, which means that basis vectors are eigenvectors of B.

OpenStudy (anonymous):

Let's say if A has totally different eigenvalues, ΛB=BΛ (Λ , the diagonal matrix of A) deduces that B is also diagonal, very well. What if Λ has the same eigenvalue, I don't think B is diagonal deduced from ΛB=BΛ. Also, AB=BA equal to ΛB=BΛ? IF A=SΛS^(-1), then SΛS^(-1)B=BSΛS^(-1), how I get ΛB=BΛ?

OpenStudy (anonymous):

I like your question. Generally speaking the statement of your theorem is not true. However, existence of common basis formed by eigenvectors of diagonalizable A and B is equivalent to AB = BA. Consider extreme case when all eigenvalues of A are the same. Then any vector of n-dimensional space Rn is an eigenvector of A. In that case the statement that A and B have common basis formed by eigenvectors is trivial. I will show proof of the general case later today.

OpenStudy (anonymous):

Your last question: "Also, AB=BA equal to ΛB=BΛ? IF A=SΛS^(-1), then SΛS^(-1)B=BSΛS^(-1), how I get ΛB=BΛ?" Answer: When A = S*DIAG*S^(-1) then B transforms to B' = S*B*S^(-1), so AB=BA => DIAG*B' = B'*DIAG. In other words, property AB = BA does not depend from choice of the basis.

OpenStudy (anonymous):

Thanks for your patient answer! There's still doubts..but I'll wait and see your proof of general case first.

OpenStudy (anonymous):

General case. 1. There exists basis formed by eigenvectors of A, since A is diagonalizable. 2. Sort the basis vectors by ascending order of corresponding eigenvalues. In that basis: a) Matrix representation of A is a diagonal matrix DIAG. b) Matrix representation of B is cellular-diagonal matrix CELL, as follows from equation AB=BA. 3. Since B-cells outside of main diagonal are zeroes, each group of basis vectors, corresponding the same eigenvalue of A, spans a B-invariant subspace. 4. Restriction of B on that subspace is also diagonalizable. 5. There exists a basis of that subspace formed by eigenvectors of B. 6. All base vectors of this basis are eigenvectors of A, since any vector of this subspace is an eigenvector of A ( as in extreme case above). 7. Union of all such base systems forms common basis in Rn, consisting of common eigenvectors of A and B.